OpenStudy (anonymous):

What is the exact distance from (–1, 4) to (6, –2) ill give medals

OpenStudy (anonymous):

7 across, 6 down. a^2 + b^2 = c^2, so 7^2 + 6^2 = c^2 49 + 36 = c^2 85 = c^2 c = 9.2195